t 1/2 +3t 1/4 +2=0 solve this problem
Found 2 solutions by pinklipgloss2007, richwmiller:
Answer by pinklipgloss2007(6)

(1/2)t+3(1/4)t+2=0
(1/2)t+(3/4)t+(2/1)=0
We have to find a common denominator in order to add the fractions.
(One fraction is better than 3)
The largest denominator we have is 4, therefore, we need to make all the other denominators 4.
What do we have to multiply by 2 to get 4? 2
What do we have to multiply by 1 to get 4? 4
But if we multiply the denominators by 4, we also have to multiply the numerators by 4!
(2/4)t+(3/4)t+(8/4)=0
(2t+3t+8)/4=0
(5t+8)/4=0
We can split this up into 2 parts.
(5t)/4 + (8/4) = 0
(5t/4)+2=0
Subtracte 2 from both sides.
(5t/4)=-2
To get the t by itself- we divide 5t/4 by 5/4. These numbers cancel out and we are left with t.
But if we divide by 5/4 on one side, we must do it on the other.
To divide a fraction by a fraction, you multiply 5/4 by it's reciprocal. That is, 4/5.
-2/1 * 4/5
t= -8/5 OR -1.6

Answer by richwmiller(17219)

t/2+3t/4=-2
2t/4+3t/4=-2
5t=-8
t=-8/5